### v2

# Better Language Models Exhibit Higher Visual Alignment

Abstract
How well do text-only Large Language Models (LLMs) naturally align with the visual world? We provide the first direct analysis by utilizing frozen text representations in a discriminative vision-language model framework and measuring zero-shot generalization on unseen classes. We find decoder-based LLMs exhibit high intrinsic visual alignment. In particular, more capable LLMs reliably demonstrate stronger generalization. Moreover, utilizing frozen LLMs leads to strong gains in cross-lingual settings, where our approach surpasses CLIP’s accuracy of with for Chinese. Our proposed method improves both robustness and generalization and also significantly reduces the need for paired data and compute, making vision-language models more accessible and adaptable. Code and models are published on our Project Page
Better Language Models Exhibit Higher Visual Alignment
Jona Ruthardt FunAI Lab University of Technology Nuremberg Gertjan J. Burghouts Intelligent Imaging TNO
Serge Belongie Department of Computer Science University of Copenhagen Yuki M. Asano FunAI Lab University of Technology Nuremberg
1 Introduction
Large Language Models (LLMs) are solely pretrained on unimodal textual data, yet are increasingly incorporated into systems that perceive and interact with the natural world (Ahn et al., 2022; Driess et al., 2023; Wayve, 2023). The lack of direct sensory experience raises fundamental questions as to what extent such models can generalize across modalities and develop a meaningful and accurate understanding of visual reality. Do these models merely regurgitate visually relevant factual knowledge from their training corpus, or do they form internal representations that correspond to real-world phenomena? Despite their successful integration into large-scale Vision-Language Models (VLMs), judging the visual capabilities already inherent to LLMs is difficult. This is not only because of widely varying training recipes and proprietary data sources but particularly due to fine-tuning with paired image-text data, which intertwines with any visual knowledge already contained in text-only models.
In contrast, Sharma et al. (2024) and Huh et al. (2024) more immediately assess the visual nature of LLMs and highlight a non-trivial degree of visual understanding and cross-modal alignment. These works compile proxy tasks or measures such as generating code to represent real-world concepts (Sharma et al., 2024) or correlating visual- with language-based representations (Huh et al., 2024). However, reliance on highly constrained and synthetic tasks with limited practical significance fails to gauge the aptitude of LLMs when deployed in more realistic settings.
To this end, we assess visual alignment—the degree to which language model representations structurally and semantically correspond to those of vision models—through the task of zero-shot open-vocabulary image classification, as popularized by CLIP (Radford et al., 2021). This involves learning a projection of language embeddings into the latent space of a vision model and selecting the label whose representation is most similar to a given image. To ensure a rigorous evaluation of true zero-shot generalization, we enforce strict disjointness between seen and unseen concepts (Lampert et al., 2009). This mitigates concept leakage, a common issue in VLMs trained on web-scale data. For example, Xu et al. (2024) find significant overlap between CLIP’s training data and common evaluation datasets and other works (Fang et al., 2022; Udandarao et al., 2024; Parashar et al., 2024; Mayilvahanan et al., 2025) demonstrate and sharp drops in recognition performance for less-frequent or truly novel concepts. In our proposed setup, generalization relies solely on the semantic information and visual knowledge encoded in language representations. By probing how well models capture visual semantics, we provide insight into their ability to structure and encode text for vision-language applications.
We test a range of language model types and find that features extracted from large modern decoder-based LLMs are more effective than classic encoder-based embeddings, like BERT. Intriguingly, we find that general LLM capability, as measured by MMLU-Pro (Hendrycks et al., 2021b), correlates positively with the model’s visual performance, as shown in Figure 1. Even off-the-shelf LLMs without embedding-specific fine-tuning exhibit strong visual representation abilities.
Finally, we integrate frozen vision and language representations into a lightweight discriminative VLM, ShareLock, demonstrating strong multimodal capabilities across a range of tasks. By capitalizing on the broad pretraining of modern LLMs, ShareLock achieves remarkable cross-lingual zero-shot generalization to non-English languages, outperforming CLIP’s performance of with for Chinese. Despite using only a fraction of the data and learnable parameters, our method approaches the performance of fully optimized CLIP models trained on orders of magnitude more paired data. The visual and linguistic expressiveness of LLM representations is particularly effective in nuanced and fine-grained tasks, resulting in above-CLIP compositional reasoning performance. Our results highlight the considerable extent to which language representations inherently capture visual structure and semantics, despite their unimodal pretraining.
Overall, the main contributions of this work are:
We systematically assess the visual alignment inherent to language models through the lens of strict zero-shot image classification.
Our analysis highlights modern LLMs as effective sources of visual knowledge, with semantically meaningful representations extractable from their internal states.
With ShareLock, we incorporate frozen LLMs with intrinsic visual alignment into VLMs, resulting in improved robustness and generalization that we demonstrate on various tasks involving classification, retrieval, multi-lingual understanding, and compositional reasoning.
2 Related Work
Visual Understanding of Large Language Models.
LLMs can infer and reason about visual content without explicit multi-modal training (Bowman, 2023). Sharma et al. (2024) tasked LLMs to draw common objects and scenes using simple shapes, indicating spatial understanding and illustrating that LLMs can conceptualize real-world settings. Various works highlight the plausibility and utility of LLM-generated descriptions of objects in the context of image classification and demonstrate that LLMs possess encyclopedic knowledge about visual characteristics (Pratt et al., 2022; Menon and Vondrick, 2023; Yang et al., 2022; Saha et al., 2024). These capabilities suggest that the extensive pretraining on large volumes of diverse textual data aids the visual understanding of LLMs. Huh et al. (2024) argue that the embedding spaces of neural networks converge towards a shared ‘platonic’ representation of reality irrespective of the concrete optimization objectives and data modality utilized during training. Similarly, we investigate the degree of visual alignment inherent to exclusively language-based representations but assess this in the practically more relevant context of zero-shot image classification and design a rigorous benchmark to measure the true generalization capabilities facilitated by such language embeddings.
Vision-Language Alignment.
Similarly to Huh et al. (2024), other recent works indicate and exploit alignment between pretrained unimodal vision and language models. Zhai et al. (2022) and Khan and Fu (2023) reveal that leveraging pretrained models and only tuning a subset of parameters can improve performance and efficiency over CLIP (Radford et al., 2021), indicating some degree of model-inherent alignment. Norelli et al. (2023) and Maniparambil et al. (2024) only rely on cross-modal correlations for training-free alignment of unimodal models. However, these studies are primarily restricted to encoder-based language models. Zhang et al. (2024) incorporate decoder-based LLMs into a CLIP-like framework but transform both vision and language representations, making it difficult to isolate the contribution of each modality during alignment. In contrast, we systematically evaluate both encoder- and decoder-based language models, examine how well their representations can be mapped to visual latent spaces, and focus on zero-shot generalization.
3 Probing LLMs for Visual Alignment
3.1 Zero-Shot Generalization
In order to quantitatively gauge the visual aptitude of language models, we draw on traditional zero-shot learning methodology (cf. Lampert et al. (2009)) and assess how their representations facilitate generalization to novel concepts. To rigorously assess true generalization performance without concept leakage from supervision with arbitrary web-scraped image-captions pairs, we split image classification datasets into seen classes (training stage) and unseen classes (testing stage) with . In the absence of image-specific captions, text-based class representations are used as supervision signals during training and for zero-shot transfer during inference. Crucially, this implies that the performance to discriminate novel concepts is contingent on the validity and cross-modal continuity of the class representations. Therefore, this setup allows us to assess the degree to which language models encode visual knowledge and semantics.
3.2 Shared Vision-Language-Locked Tuning
To map textual inputs into visual latent spaces, we draw inspiration from late-fusion architectures in CLIP-like models. Texts are first encoded using a language model and subsequently projected into the -dimensional latent space of the vision encoder via a learnable projection network . The latent representation for a given input image or caption is therefore computed by and , respectively. After normalization, their similarity is computed as the cosine similarity, given by the dot product of the embeddings.
During training, only the lightweight projection network is optimized, while the pretrained vision and language backbones remain frozen. A contrastive loss encourages alignment by pulling textual representations closer to their corresponding image embeddings while pushing them away from non-matching ones, as in (Radford et al., 2021). For an image-text pair in a batch with items, it is given by
for both alternated modalities pairings and with being a fixed temperature parameter. Given a set of classes and their corresponding textual class representations , the predicted class for a sample is obtained via
3.3 Experimental Setup
Datasets.
For a comprehensive evaluation, we select four datasets: AWA2 (Xian et al., 2017), CUB (Wah et al., 2011), FGVCAircraft (Maji et al., 2013), and ImageNet+, spanning natural and human-made artifacts, coarse and fine-grained categories, and varying scales (). ImageNet+ considers ImageNet-1k (Deng et al., 2009) classes as seen concepts and uses the 500 most populated ImageNet-21k classes as unseen ones. For AWA2 and CUB, we use splits by Xian et al. (2017) while randomly dividing aircraft types into 50 seen and 20 unseen classes. We report the average per-class classification accuracy across the datasets as a measure of visual generalization ability facilitated by the language embeddings.
Pretrained Unimodal Backbones.
Given its strong performance, broad pretraining regime, and popularity, the ViT-L/14 variant of the self-supervised DINOv2 model family (Oquab et al., 2023) is the default vision backbone. Global image embeddings are obtained through the CLS token. Language representations are extracted from encoder-based models through mean token pooling or directly via the CLS token if fine-tuned on sentence-level representation tasks. For decoder-based LLMs other than NV-Embed (Lee et al., 2024), features are extracted through last-token pooling (details in Appendix). Frozen vision and language model features are initially precomputed and stored for direct re-use in subsequent epochs.
Training.
Given the data-constrained setting, we optimize a linear projection network using Adam (Kingma and Ba, 2014) with a learning rate following a cosine schedule with a maximum value of . Gradient clipping to a global norm of and weight decay of are applied. The loss of Eqn. 1 is applied with , and models are trained until convergence on a randomly chosen validation split or for a maximum of k steps with a batch size of and five different initialization seeds. Dropout (Srivastava et al., 2014) with is applied during training.
4 Language-driven Visual Generalization
Utilizing the zero-shot evaluation methodology outlined, we investigate critical factors that promote the generalization of language representations to illuminate language models’ visual capabilities.
LLM representations encode visual knowledge.
The class-wise supervision and limited concept diversity of conventional image classification datasets can impede vision-language alignment but permit more sophisticated semantic class representations beyond simple template-based targets as typically used with CLIP-like models (Radford et al., 2021; LAION AI, 2022). We thus examine how the nature and information content of different textual class representations111Details about the characteristics and acquisition of these class representations are elaborated in Section B impact generalization performance. The results are summarized in Table 1.
| Class | LLM | original | ||
| Language Model | Names | Wikip. | Wikip. | |
| Enc. | BERT-Large (Devlin et al., 2019) | |||
| ModernBERT (Nussbaum et al., 2024) | ||||
| all-roberta-v1 (Liu et al., 2019) | ||||
| T5-XL (Raffel et al., 2020) | ||||
| SentenceT5-XXL (Ni et al., 2022) | ||||
| Flan-UL2 (10B) (Tay, 2024) | ||||
| Dec. | Gemma-2 9B (Mesnard et al., 2024) | |||
| Llama-3 8B (Dubey et al., 2024) | ||||
| NV-Embed-v2 (Lee et al., 2024) |
In line with previous studies (Pratt et al., 2022; Menon and Vondrick, 2023; Yang et al., 2022; Saha et al., 2024), we find that the addition of auxiliary information, such as Wikipedia articles, results in improved performance for most language models. We also find that LLM-generated articles describing a class in the style of Wikipedia (“LLM Wikip.” in Table 1) can provide strong targets during multi-modal alignment, achieving the best overall performance of . Interestingly, relying on strictly human-curated data in the form of actual Wikipedia articles tends only to provide marginal benefits, for example, from and , for NV-Embed and Llama-3. Thus, LLMs can effectively absorb and interpolate substantial amounts of factual information from their training data, positioning them as valuable sources of visually relevant knowledge.
Decoders excel in visual concept representation.
A new insight resulting from this analysis is the competitiveness of decoder-based language models for representing visual concepts. Compared to encoders, we find that representing inputs with decoders can result in higher performance for visual tasks, mirroring a recently emerging trend in the language domain (Lee et al., 2024; Springer et al., 2024). This is likely because these LLMs are trained with vast (and proprietary) training data and are typically much larger than common encoder LMs. In particular, Gemma-2 9B (Mesnard et al., 2024) emerges as the best performer across various types of input data with a maximum performance of . The best-performing encoder, Flan-UL2 (Tay, 2024), only achieves a score of on the same input data, with T5- and particularly BERT-based models exhibiting significantly lower performances. In contrast, other LLMs like Llama or NV-Embed mirror Gemma-2’s performance level more closely.
LLM and its visual performance are correlated.
In Figure 1, we compare various LLMs by the visual generalization ability they possess, as well as their MMLU-Pro (Wang et al., 2024b) score taken from the Open LLM Leaderboard (Fourrier et al., 2024), a common metric to measure LLM performance. We find that the general capability of language models is strongly correlated with their ability to perform well on the visual tasks (Pearson coefficient : ). Within and across model families, we see improved visual generalization as the capacity and capabilities of models increase. Since models steadily improve in the language domain, our evaluation protocol will be helpful in assessing whether the trend of increasing visual understanding will continue in future LLM models. If this holds, VLMs that incorporate LLMs can piggyback off and benefit from developments in the language modeling domain.
| Model | Dataset | Size | IN-1k | IN-R | IN-A | IN-S | Pets | Flowers | Cars | Aircraft | Avg |
| LiT | COCO | 83k | |||||||||
| ASIF | COCO | 83k | |||||||||
| ShareLock | COCO | 83k | |||||||||
| LiT | CC3M Subset | 563k | |||||||||
| ASIF | CC3M Subset | 563k | |||||||||
| ShareLock | CC3M Subset | 563k | |||||||||
| CLIP | CC3M | 2.8M | |||||||||
| SLIP | CC3M | 2.8M | |||||||||
| LaCLIP | CC3M | 2.8M | |||||||||
| LiT | CC3M | 2.8M | |||||||||
| ShareLock | CC3M | 2.8M | |||||||||
| DataComp | CPool-S | 3.84M | |||||||||
| CLIP | CC12M | 12M | |||||||||
| SLIP | CC12M | 12M | |||||||||
| LaCLIP | CC12M | 12M | |||||||||
| LiT | CC12M | 8.5M | |||||||||
| ShareLock | CC12M | 8.5M | |||||||||
| DataComp | CPool-M | 38.4M | |||||||||
| DataComp | CPool-L | 384M | |||||||||
| CLIP | Proprietary | 400M |
5 LLMs for General-Purpose VLMs
The previous section revealed the intriguing capabilities of decoder-based LLMs to encode visually relevant information in their representations. Next, we investigate whether this inherent visual alignment can be leveraged to obtain effective and efficient multimodal models by integrating off-the-shelve LLMs directly into CLIP-like VLMs. For this, we relax the strict zero-shot setup and utilize larger-scale image-caption datasets to explore the benefits and limitations of general-purpose VLMs that fuse existing frozen foundation models with minimal data and compute.
5.1 Experimental Setup
Methodology.
We propose integrating LLMs into a CLIP-like framework through “Shared Vision-Language-Locked Tuning“ (ShareLock), extending the architecture and training setup introduced in Section 3. As illustrated in Figure 2, this design places the LLM at the core of the model and necessitates expressive language representations for strong VLM performance. Given the increased availability of training data compared to the strict zero-shot setting, we increase the projection network to four layers with a hidden dimension of . The additional parameters, combined with ReLU activations between layers, enhance ShareLock’s ability to capture complex, non-linear cross-modal correspondences. Additionally, the maximum number of steps is increased to 5000.
Datasets.
Our investigation focuses on leveraging LLMs with minimal additional paired data and explores how unimodal embeddings can drive robust multimodal performance with minimal supervision and alignment. As a result, our evaluation is limited to comparably small paired datasets.
COCO Captions. COCO Captions (Chen et al., 2015) contains 83k images with multiple human-written captions per image, from which we randomly sample during training. CC3M. Conceptual Captions (Sharma et al., 2018) comprises 2.8M filtered web-scraped image-alt-text pairs. We also utilize a smaller subset with more balanced concept coverage designed for LLaVA (Liu et al., 2023). CC12M. CC12M (Changpinyo et al., 2021) expands CC3M by 12M image-text pairs, enabling evaluation at larger scales. Due to expired links, our version contains about 8.5M samples.
Training speed and storage.
Precomputing features for the CC3M-Llava subset (563k pairs) using Llama-3 8B requires around 8 hours on a single A100 40GB GPU. Extracting DINOv2 features and optimizing the MLP-based projection network take 1 GPU each, totaling roughly 10 GPU hours. This is paired with a significant storage reduction from over 80GB for the raw data to just 12GB for precomputed features.
Evaluation.
We employ a comprehensive suite of evaluations to assess ShareLock’s capabilities across a wide range of tasks. Based on the publicly available CLIP Benchmark (LAION AI, 2022), we gauge the models’ zero-shot classification abilities across diverse datasets: ImageNet-1k (Deng et al., 2009), ImageNet-R(endition) (Hendrycks et al., 2021a), ImageNet-A(dversarial) (Hendrycks et al., 2021c), ImageNet-S(ketch) (Wang et al., 2019), Oxford-IIIT Pets (Parkhi et al., 2012), Oxford Flowers (Nilsback and Zisserman, 2008), Stanford Cars (Krause et al., 2013), and FGVCAircraft (Maji et al., 2013). We also provide qualitative text-to-image retrieval results on ImageNet for CC3M-trained models. Moreover, the challenging compositionality Winoground task (Thrush et al., 2022) is explored.
Baselines and comparisons.
We compare our straightforward and economical way of incorporating LLMs into VLMs to more conventional CLIP-like methods, particularly emphasizing data-efficient alignment approaches. Alongside the original ViT-B/16 variant of CLIP (Radford et al., 2021), we test against several CLIP-like models trained on public datasets of different scales and modified learning objectives (Fan et al., 2023; Mu et al., 2022; Gadre et al., 2023; Mu et al., 2022). Leveraging pretrained models, we evaluate how ShareLock compares to LiT (Zhai et al., 2022) and ASIF (Norelli et al., 2023) by reproducing these methods on smaller-scale datasets. For LiT baselines, we initialize the language encoder with pretrained BERT weights (Devlin et al., 2019), following Zhai et al. (2022). When comparing LiT, ASIF, and ShareLock models, the exact same precomputed input features (barring the language component of LiT) are used.
5.2 Comparison to Conventional VLMs
Classification.
LLMs can directly be leveraged profitably in vision-centric tasks and outperform conventional models trained on similar small-scale datasets as demonstrated by the IN-1k accuracies in Table 2. ShareLock’s accuracy on CC3M substantially exceeds both CLIP () and LiT (), despite the latter using the same vision features and fully fine-tuning the language component. Even with larger datasets like CC12M, where full fine-tuning becomes more viable, minimal transformations on top of LLM representations maintain a competitive advantage of over LiT and CLIP. Compared to the training-free ASIF, optimizing a small number of parameters proves advantageous and forgoes the reliance on large and diverse reference datasets and extensive compute during inference.
Beyond competitive performance on general-purpose classification, leveraging strong representations from pretrained models enables increased robustness to out-of-distribution image inputs as seen in columns “IN-R“ to “IN-A“ of Table 2, surpassing the robustness of the original CLIP model despite being exposed to a fraction of the training data (8.5M vs. 400M).
The fine-grained nature of certain classification problems (cols. “Pet“ to “Aircraft“ in Tab. 2) demands larger-scale datasets with more diverse and nuanced concepts included as minute visual differences may be insufficiently captured in the text space. Consequently, low performance on small-scale datasets can be observed across all methods. Nonetheless, the LLM representations still contain visually valuable signals, enabling ShareLock to surpass alternative methods trained on the same data in 15/16 cases and demonstrating effective utilization of intrinsic LLM knowledge to generalize to truly novel concepts.
Multi-Lingual Understanding.
Most popular multimodal datasets consist primarily of English captions. Consequently, VLMs like CLIP and LiT experience significant performance drops when performing inference in other languages, as shown in Table 3 on multi-lingual ImageNet LAION AI (2022). In contrast, the broader and typically multi-lingual pretraining of LLMs enables ShareLock to harness cross-linguistic image-text consistencies, greatly mitigating the performance loss in non-English languages. Even with substantially fewer training samples, ShareLock surpasses the original CLIP model, achieving an accuracy of versus for Chinese and versus for Italian. In comparison, both LiT and CLIP models similarly trained on CC12M demonstrate near-random performance in these languages. This transfer of capabilities makes the inclusion of LLMs especially attractive for low-resource languages with little available or high-quality multimodal data.
| Model | Dataset | [Size] | EN | CN | JP | IT |
| LiT | COCO | 83k | ||||
| ShareLock | COCO | 83k | ||||
| CLIP | CC12M | 12M | ||||
| LiT | CC12M | 8.5M | ||||
| ShareLock | CC12M | 8.5M | ||||
| DataComp | CPool-M | 38.4M | ||||
| DataComp | CPool-L | 384M | ||||
| CLIP | Proprietary | 400M |
Compositionality.
Late-fusion VLMs often struggle with nuanced textual and fine-grained compositional differences, as seen in benchmarks like Winoground (Thrush et al., 2022) and SugarCrepe (Hsieh et al., 2023). Despite the intriguing linguistic and generative abilities of LLMs, their representations fail to adequately reflect fine linguistic differences in the vision-language contrastive setting. While ShareLock improves image scores over CLIP ( vs. ) and thus more reliably selects the correct image given a textual description, it still falls short of significant above-random performance and remains far from human-level capability. However, the low performance on compositionality tasks might partly be an architectural limitation, as recent works (Zhang et al., 2024; Jose et al., 2024) have indicated limitations of solely aligning language representation to the vision space, as suggested by Zhai et al. (2022). Thus, also applying transformations on top of vision features can be beneficial, especially in retrieval and detail-oriented settings.
| Model | Dataset | [Size] | Text | Image | Group |
| Human | |||||
| Chance | |||||
| LiT | COCO | 83k | |||
| ASIF | COCO | 83k | |||
| ShareLock | COCO | 83k | |||
| CLIP | CC12M | 12M | |||
| LiT | CC12M | 8.5M | |||
| ShareLock | CC12M | 8.5M | |||
| DataComp | CPool-M | 38.4M | |||
| DataComp | CPool-L | 384M | |||
| CLIP | Propriatary | 400M |
Data scaling.
Figure 4 illustrates that ShareLock achieves comparable performance to CLIP and DataComp models using orders of magnitude less data. Utilizing frozen LLM features is especially effective in low-data regimes, consistently outperforming conventional CLIP-like models, further underlining their visually-relevant semantic content and capacity to facilitate generalization.
5.3 Qualitative Results
In addition to quantitative evaluations, Figure 3 demonstrates ShareLock’s strong text-image alignment across diverse prompts, showing particular advantages over CC3M-trained CLIP and LiT models for both fine-grained (e.g., “a photo of a BMW") and abstract (e.g., “[…] heavy seas") queries.
5.4 Choice of Language Model
As the nature and quality of the frozen language features are of great significance in the proposed architecture, we ablate the choice of language model on the CC3M dataset. Reflecting the insights from our investigation in Section 4, Table 5 highlights the potential of decoder-based models for vision-language tasks. Although BERT encoders serve as the starting point in LiT models, they perform poorly without fine-tuning. Similarly, all-roberta-v1 (Liu et al., 2019) improves significantly but remains inferior to LLM-based representations, despite being highlighted by Maniparambil et al. (2024) for its high inherent visual alignment. In contrast, frozen decoder-based representations consistently surpass BERT-based ones, with gains of to , showcasing the richness of strong LLM representations from text-only pretraining. Even naive last-token-pooling is suitable for extracting visual information from LLMs and the more sophisticated multi-token embedding approach of NV-Embed (Lee et al., 2024) fails to materialize in consistent and notable performance gains over off-the-shelf Llama and Gemma models.
Further analysis into the role of vision backbones and using more sophisticated projection network architectures can be found in Appendix D.
| Type | Language Model | IN1k | IN-R | IN-A | Pets | Cars | Aircraft |
| Enc. | BERT-Base | ||||||
| all-roberta-v1 | |||||||
| Dec. | Llama-3 8B | ||||||
| Gemma-2 9B | |||||||
| NV-Embed-v2 |
6 Conclusion
We systematically investigate the visual capabilities and inherent alignment of unimodal language models. Our analysis demonstrates that general LLM quality, as measured by MMLU-Pro, correlates with visual aptitude and that decoder-based models effectively generalize across modalities. By integrating off-the-shelf LLMs into a lightweight CLIP-like architecture, we leverage their large-scale pretraining, intrinsic knowledge of the visual world, and multilingual capabilities, achieving competitive performance with conventional VLMs trained on significantly larger datasets. Our findings offer a deeper understanding of state-of-the-art language models and highlight their potential for broader adoption in vision-centric applications.
7 Limitations
While we compare a wide range of publicly available encoder- and decoder-based language models, attributing concrete performance differences on mainly architectural differences is not possible due to the different pretraining objectives, training datasets, and number of parameters used in these models. As the primary focus of this work is on highlighting the visual alignment inherent to unimodal language models and their incorporation into data- and compute-efficient vision-language models, our investigation is limited to datasets with up to 12M image-caption pairs. Scaling ShareLock to larger datasets for further performance improvements is left for future work. Although pre-computing features only once is possible due to locked vision and language backbones, using LLMs with billions of parameters significantly increases computational costs of forward passes compared to conventional encoder architectures.
8 Acknowledgments
We gratefully acknowledge the ELLIS Unit Amsterdam for providing funding for a research visit to Copenhagen. We thank SURF for providing GPU cluster access and support during this project. We acknowledge the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) with hardware funded by the German Research Foundation (DFG). This work was supported in part by the Pioneer Centre for AI, DNRF grant number P1.
Appendix A Reproducibility Statement
We acknowledge and emphasize the importance of reproducibility in our work and take active measures to facilitate reproducibility efforts. Besides providing comprehensive documentation of our methods throughout the main paper, with additional details in the supplementary materials, we will publish source code for the proposed ShareLock model.
Our use of existing models aligns with their intended purpose and is carried out in an academic setting. Any data used follows its original access conditions, and all artifacts produced are intended for research purposes only. Only publicly available resources and scientific artifacts are used.
Appendix B Acquisition of Textual Class Representations
More details about the characteristics and acquisition of these class representations are provided in Section B of the appendix. Besides the template-based targets proposed by Radford et al. (2021) that solely substitute the respective class names, we generate more comprehensive auxiliary information about classes (e.g., Wikipedia-style articles) using the instruction-tuned version of the Llama-3-8B model and acquire human-curated information from Wikipedia (details provided in B).
Class representations are essential for facilitating the knowledge transfer between classes in the traditional definition of zero-shot learning. Compared to attributes or other forms of class semantics, language-based class representations are more conveniently accessible at various scales and may come in diverse manifestations. The advent of LLMs adds further possibilities for generating and obtaining such auxiliary information. The following paragraphs specify the respective properties and acquisition process. Here, all LLM-based class representations are generated using the instruct-tuned version of LLama-3 8B.
Class Names.
A set of 80 human-engineered prompt templates in the style of "a photo of a <class name>" are adopted from Radford et al. (2021).
Wikipedia Page.
Being a comprehensive and mostly factually correct source of information, Wikipedia constitutes an interesting source of auxiliary information in the context of zero-shot classification. To obtain class-article correspondences, class names are automatically matched with page names, after which additional manual quality checks are performed. Nonetheless, an ideal match does not always exist due to high class specificity or generality, in which case superordinate articles are considered or template-based fallbacks are employed.
LLM-based Wikipedia Style Articles.
Despite being specifically prompted for articles mimicking Wikipedia, the Llama-3-generated texts tend to show significant differences in style compared to their real counterparts.
As the lengthy nature of Wikipedia(-style) articles might dilute the information content captured by the language embeddings, the texts are split into individual sentences, which are used as targets during training. For all types of class representations, predictions are made by aggregating class scores through averaging over all individual class-specific texts.
Appendix C In-Depth Analysis of Visual LLM Generalization
As outlined in Section 4, we find that the general capability of language models is strongly correlated with their ability to perform well on visual tasks. While this is also true for members of the Phi-3 family of models, their absolute visual generalization scores are notably lower compared to models of similar size and capability, as seen in Figure 5. This discrepancy likely illustrates the effects of the extensive data curation and synthetic data creation utilized in Phi-3, which might remove visual information to favor tokens that promote reasoning abilities. Thus, a lack of exposure to sufficient factual knowledge about real-world conditions may impede the formation of visually informed representations.
As seen in Figure 5 and Table 12, the text encoder taken from the CLIP ViT-B/16 model constitutes a strong baseline that is not yet reached by any of the current state-of-the-art LLMs. However, considering the explicit vision alignment on 400M image-text pairs makes the strong generalization abilities unsurprising. On the contrary, obtaining a score of as the best performing unimodal model, Gemma-2 9B (Mesnard et al., 2024) is close to CLIP’s despite no multimodal exposure. This further underscores the remarkable degree of visual alignment inherent to decoder-based language models.
LLMs are often subject to additional task-specific fine-tuning. Table 6 compares different models and their derivatives tuned on instruction and multimodal data. While vision-tuned variants are available for Phi-3 (Microsoft, 2024), Qwen2 (Wang et al., 2024a), and Vicuna v1.5 (Zheng et al., 2023), XTuner’s LLaVA model (XTuner Contributors, 2023) constitutes the source for the visual Llama-3 variant. Across all models, with the exception of Llama-3, the impact of fine-tuning is minor, typically shifting performance by only a few decimal points in the case of Phi-3 and Vicuna. Interestingly, Llama-3’s performance declines significantly after fine-tuning ( and for instruction and visual tuning), contrasting with the generally stable results of other models. Neither instruction-based nor visual fine-tuning shows a clear and consistent advantage in improving overall performance. These insights are also reflected in Figure 5 and Table 12. Ultimately, the base model’s architecture and training regime are more significant in determining performance than post-hoc fine-tuning strategies.
| Llama-3 8B | Phi-3 Mini | Qwen2 7B | Llama-2/Vicuna 7B | |
| None | 44.4 | n/a | 40.2 | 42.5 |
| Instruct | 42.7 | 36.9 | 41.8 | 42.3 |
| Visual | 34.0 | 37.0 | 40.3 | 42.6 |
Appendix D Additional Ablations
D.1 Projection Network Architecture
The multi-layer perceptron (MLP) projection networks of ShareLock as introduced in Section 3 are conceivably simple. As these are the only unfrozen and tunable parts of the model architecture and thus responsible for aligning vision and language inputs, they are of particular significance to aptly process and transform the inputs. Following Zhai et al. (2022), no transformation to the vision inputs is applied for any of the architectures. With a hidden size of 4096 and four layers, the MLP processing the language features comprises approximately 53M parameters.
In addition to the straightforward MLP-based networks, also more sophisticated Transformer-based architectures are inspired by recent works. First introduced as part of the BLIP-2 model (Li et al., 2023), the Q-Former is a lightweight Transformer-based model that extracts features from an input modality using cross-attention with learnable query tokens. Similarly, albeit introduced in a different context, NV-Embed (Lee et al., 2024) uses a latent attention layer to pool language tokens and receive a global embedding. Slight adjustments are made to both baseline architectures to better suit late-fusion vision-language modeling, which we will denote as Q-Transformer and NV-Transformer. While features are extracted via last-token-pooling (see Figure 6) when used with MLP projection networks, all tokens of the input sequence are considered for the alternative architectures. The hyperparameters were selected based on the implementation details suggested in the original publications and to approximately match the MLP baseline in learnable parameter count. Both the Q-Transformer and the NV-Transformer projection networks have a token dimension of 1024 in the Transformer parts of the models, eight learnable queries (Q-Transformer), and key/values (NV-Transformer). Whereas the Q-Transformer consists of 3 blocks and 4 attention heads, NV-Transformer comprises a total of four layers with eight cross-attention heads each.
The choice of projection network architecture is ablated in Table 7. The evaluated models use DINOv2-ViT-B/14 as their vision backbone. While no single architecture consistently scores best, the MLP-based ShareLock configuration performs competitively compared to NV-Transformer and Q-Transformer throughout the evaluation datasets. Additionally, Transformer-based architectures entail increased computational complexity due to the more evolved attention mechanism and processing of more tokens, making MLPs an attractive choice from an efficiency perspective as well. These results suggest that the additional information contained across all tokens of an input is not significantly more adjuvant compared to solely considering the last token representation as is done with the MLP.
| Architecture | IN1k | IN-R | IN-A | Pets | Cars | ESAT |
| NV-Transformer | ||||||
| Q-Transformer | ||||||
| MLP |
D.2 Loss Function
The use of the Sigmoid Loss (SigLIP) has unlocked further efficiency and performance enhancements in the standard CLIP regime (Zhai et al., 2023). However, no substantial gains are found when using SigLIP in the ShareLock framework as presented in Table 8.
| Loss Function | IN1k | IN-R | IN-A | Pets | Cars | ESAT |
| SigLIP Loss | ||||||
| CLIP Loss |
D.3 Choice of Vision Encoder
| Language Model | IN1k | IN-R | IN-A | Pets | Cars | Aircraft |
| ResNet101 (sup.) | ||||||
| ConvNextV2 (sup.) | ||||||
| ViT-L (sup.) | ||||||
| DINOv2-L (self-sup.) |
As ShareLock is agnostic to the utilized vision encoder, we ablate variants with differing architectures and supervision regimes in Table 9. Since language embeddings map to the vision space, encoder choice is a crucial factor, as reflected in notable performance differences. Unlike DINOv2, the other models were only trained on ImageNet and exhibit lower robustness and generality, highlighting the benefits of broad pretraining across diverse concepts – even without explicit supervision.
One advantage of CLIP is its favorable scaling characteristics when increasing the size of the vision encoder (Radford et al., 2021). To validate if comparable trends are present in ShareLock, we vary DINOv2-based backbones ranging from Small to Giant vision transformers and present the results in Table 10. Indeed, ShareLock also profits from scaling up the vision backbones with the average scores increasing by and , when moving from the Small to the Base and from the Base to the Large DINOv2 models, respectively. However, the benefits of scale start to level off thereafter, and only marginal differences are present when utilizing representations from the Giant vision encoder.
| Variant | IN1k | IN-R | IN-A | Pets | Cars | ESAT |
| Small | ||||||
| Base | ||||||
| Large | ||||||
| Giant |
Appendix E Extended Visual Generalization Results
Due to limited space, the reported results on visual generalization in Table 1 and Figure 1 of the main paper are averaged accuracies across five seeds and four datasets (AWA2 (Xian et al., 2017), CUB (Wah et al., 2011), FGVCAircraft (Maji et al., 2013), and ImageNet+). For increased transparency, the results are presented without dataset-level aggregation in Tables 11 and 12.
| Class Representation | Type | Language Model | AWA2 | CUB | Aircraft | ImageNet+ | Average |
| Class Names | Decoder | Llama-3 8B | |||||
| Llama-3.1 8B | |||||||
| Gemma 7B | |||||||
| Gemma-2 9b | |||||||
| NV-Embed-v2 | |||||||
| Encoder | T5-XL | ||||||
| BERT-Large | |||||||
| Flan-UL2 | |||||||
| ModernBERT-embed | |||||||
| sentence-t5-xxl | |||||||
| All-roberta-v1 | |||||||
| Pseudo Wikipedia Page | Decoder | Llama-3 8B | |||||
| Llama-3.1 8B | |||||||
| Gemma 7B | |||||||
| Gemma-2 9b | |||||||
| NV-Embed-v2 | |||||||
| Encoder | T5-XL | ||||||
| BERT-Large | |||||||
| Flan-UL2 | |||||||
| ModernBERT-embed | |||||||
| sentence-t5-xxl | |||||||
| All-roberta-v1 | |||||||
| Wikipedia Sentences | Decoder | Llama-3 8B | |||||
| Llama-3.1 8B | |||||||
| Gemma 7B | |||||||
| Gemma-2 9b | |||||||
| NV-Embed-v2 | |||||||
| Encoder | T5-XL | ||||||
| BERT-Large | |||||||
| Flan-UL2 | |||||||
| ModernBERT-embed | |||||||
| SentenceT5-XXL | |||||||
| All-roberta-v1 |
| Language Model | AWA2 | CUB | Aircraft | IN+ | Avg |
| CLIP-B/16 Text Encoder | |||||
| Qwen2-0.5B | |||||
| Qwen2-0.5B-Instruct | |||||
| Qwen2-1.5B | |||||
| Qwen2-1.5B-Instruct | |||||
| Qwen2-7B | |||||
| Qwen2-7B-Instruct | |||||
| Qwen2.5-0.5B | |||||
| Qwen2.5-0.5B-Instruct | |||||
| Qwen2.5-1.5B | |||||
| Qwen2.5-1.5B-Instruct | |||||
| Qwen2.5-14B | |||||
| Qwen2.5-14B-Instruct | |||||
| Qwen2.5-32B | |||||
| Qwen2.5-32B-Instruct | |||||
| Qwen2.5-3B | |||||
| Qwen2.5-3B-Instruct | |||||
| Qwen2.5-7B | |||||
| Qwen2.5-7B-Instruct | |||||
| gemma-2-2b | |||||
| gemma-2-2b-it | |||||
| gemma-2-9b | |||||
| gemma-2-9b-it | |||||
| Llama-3.1-8B | |||||
| Llama-3.2-1B | |||||
| Llama-3.2-1B-Instruct | |||||
| Llama-3.2-3B | |||||
| Llama-3.2-3B-Instruct | |||||
| Meta-Llama-3-8B-Instruct | |||||
| Phi-3-medium-128k-instruct | |||||
| Phi-3-medium-4k-instruct | |||||
| Phi-3-mini-128k-instruct | |||||
| Phi-3-mini-4k-instruct | |||||
| Phi-3.5-mini-instruct | |||||
| Falcon3-10B-Base | |||||
| Falcon3-10B-Instruct | |||||
| Falcon3-1B-Base | |||||
| Falcon3-3B-Base | |||||
| Falcon3-3B-Instruct | |||||
| Falcon3-7B-Base |
Appendix F Supplementary Qualitative Results
Figure 7 provides additional qualitative insights into the retrieval ability of CLIP, LiT, and ShareLock models trained on CC3M. ShareLock demonstrates visual understanding across a wide array of domains and levels of abstraction.